### v1

# A Diffusion Model to Shrink Proteins While Maintaining Their Function

Abstract
Many proteins useful in modern medicine or bioengineering are challenging to make in the lab, fuse with other proteins in cells, or deliver to tissues in the body, because their sequences are too long. Shortening these sequences typically involves costly, time-consuming experimental campaigns. Ideally, we could instead use modern models of massive databases of sequences from nature to learn how to propose shrunken proteins that resemble sequences found in nature. Unfortunately, these models struggle to efficiently search the combinatorial space of all deletions, and are not trained with inductive biases to learn how to delete. To address this gap, we propose SCISOR, a novel discrete diffusion model that deletes letters from sequences to generate protein samples that resemble those found in nature. To do so, SCISOR trains a de-noiser to reverse a forward noising process that adds random insertions to natural sequences. As a generative model, SCISOR fits evolutionary sequence data competitively with previous large models. In evaluation, SCISOR achieves state-of-the-art predictions of the functional effects of deletions on ProteinGym. Finally, we use the SCISOR de-noiser to shrink long protein sequences, and show that its suggested deletions result in significantly more realistic proteins and more often preserve functional motifs than previous models of evolutionary sequences.
1 Introduction
As protein design becomes easier, more protein constructs are built for bioengineering, more protein medicines are being packaged for delivery to particular tissues, and, of course, more protein is being synthesized in the lab. Unfortunately, many important proteins are challenging to make, engineer, and deliver, due to their long sequences. Methods to build shorter versions of these proteins are expensive and often only narrowly applicable. Typically, experimentalists look for shorter homologues, which may not exist, and put them through costly optimization campaigns [Huang2022-hl]. Or, for proteins which function by well-characterized, simple biophysical interactions, experimentalists shrink sequences by running costly and extensive physical simulations [Zhao2023-uf].
Ideally we could instead learn how to shrink proteins using models trained on databases of protein sequences in nature – these models learn the constraints evolution has put on sequences across life and could shrink proteins to avoid breaking their function. Unfortunately, these large models [Notin2022-qw, Nijkamp2022-ip] struggle to effectively search through the massive space of all possible shrunken versions of a protein. They may also lack the inductive bias to predict the effect of deletions, having not been explicitly trained to do so. In principle, the first issue could be solved by diffusion models of protein sequences, like EvoDiff, which are effectively trained to plan series of many mutations and end with sequences that resemble those found in nature [Alamdari2023-nj, Luo2022-ha]. However, current diffusion frameworks can only train models that perform substitution mutations – they cannot suggest deletions.
We propose a new diffusion model of evolutionary sequences that learns to shorten sequences — Sequence Contraction with InSertion-Only noising pRocess (SCISOR). SCISOR adds noise to natural sequences by inserting random letters until they effectively become long random sequences; then it train a de-noiser to reverse this process by planning deletions that result in sequences that resemble those found in nature (Fig. 1a). In summary:
  • We introduce SCISOR, a new discrete diffusion framework that trains a de-noiser to generate sequences by learning to delete.
  • We solve practical limitations of previous discrete diffusion models, enabling us to scale SCISOR up to large-scale evolutionary sequence data.
  • We show that among large-scale diffusion models, SCISOR achieves competitive model fit for protein sequences.
  • We show that the inductive biases of SCISOR allows it to make state-of-the-art predictions of the effects of deletions on protein functions in the lab in ProteinGym.
  • Finally, we show that SCISOR shortens proteins while better maintaining their structure and functional motifs than methods using previous models of protein sequences.
We release our code at https://github.com/baronet2/SCISOR and model weights for SCISOR small (35M), medium (150M), and large (650M) trained on UniRef50 and UniRef90 at https://huggingface.co/SCISOR/SCISOR.
Figure 1: SCISOR is a diffusion model trained to make deletions that arrive at a natural protein sequence. We can use SCISOR to shrink proteins while maintaining their function. (a) We add random insertions to protein sequences from nature and train SCISOR to reverse these insertions. (b) Applying SCISOR diffusion to natural proteins, we get smaller proteins that are predicted to preserve parts of the tertiary structures of the original sequence. We show SCISOR samples of Q8NFU3 at 0, 5, 10, 20, and 50% deletion with structures predicted by OmegaFold [Wu2022-ma].
2 Background
Say we have a protein sequence $X$ made up of $L$ letters $X^{(1)}X^{(2)}\cdots X^{(L)}$ belonging to the alphabet of 20 amino acids $\mathcal{B}$. Our goal is to remove $M$ letters from $X$ to make a $\tilde{X}=X^{(j_{1})}X^{(j_{2})}\cdots X^{(j_{L-M})}$ with $j_{1}<j_{2}<\dots,j_{L-M}$, that is still functional. Most random sets of deletions degrade the function of the protein, so we need to predict which deletions are unlikely to break the protein. Unfortunately there is very little data of sequence, shrunk-sequence pairs (X,X~)($\tilde{X}$) to learn from; we must instead learn to predict functional shrunk proteins using other available data.
Models of evolutionary sequences   One way we can learn how to shrink proteins is by learning from modern huge datasets of natural proteins. Indeed we can attempt to learn what a natural protein looks like in these databases; then we can pick a shrunken protein $\tilde{X}$ so that it looks natural and is therefore likely to be functional [^1] Note this does not guarantee our goal that $\tilde{X}$ have the same function as $X$. But if two functional proteins have similar sequences then they often have related function [Mistry2013-ky]. See also Sec. 9.. In practice, we can train huge generative models to generate natural proteins and use their likelihoods as a measure of naturalness [Riesselman2018-nm, Rives2021-co, Notin2022-qw, Nijkamp2022-ip, Lin2022-by]. Indeed, these likelihoods have been shown to be accurate predictors of whether single-letter-deletions will harm the function of a protein [Notin2022-qw].
Unfortunately, the models that are typically used to fit this data, such as BERT-style [Rives2021-co, Lin2022-by] and autoregressive models [Notin2022-qw, Nijkamp2022-ip], struggle to search over the combinatorial space of all $\binom{L}{M}$ possible sets of deletions to find an ideal $\tilde{X}$. Ideally, we would have a model that can plan a number of deletions that arrive at a functional protein sequence. We also speculate that a model that learns directly how to delete would make more accurate predictions and designs.
Discrete diffusion   To effectively search through a large mutational space, we could model the data with discrete diffusion models. These models generate samples by starting with a random sequence and applying mutations to arrive at a realistic sequence. In particular, a sequence is sampled from a simple distribution $X_{1}\sim q(X_{1})$ and then it is transformed from time $t=1$ to $t=0$ using a de-noiser $q_{\theta}((X_{t})_{t=0}^{1}\mid X_{1})$ so that $X_{0}$ looks like a sequence from the data generating distribution [campbell2022continuous].
Diffusion models can therefore be used to search for sets of many mutations to a sequence, $X$, which result in a realistic looking sequence. To do so, one sets $X_{s}=X$ for some $s$ and then de-noises using the diffusion model by sampling a path $q((X_{t})_{t=0}^{s}|X_{s})$, giving a realistic $X_{0}$ near $X_{s}$. Indeed, this procedure has been used to suggest mutations to optimize sequences [Luo2022-ha, Gruver2023-sf].
To train a de-noiser $q_{\theta}$, we first define a forward process $p((X_{t})_{t=0}^{1})$ which takes samples from our target distribution $X_{0}\sim p(X_{0})$ and applies random noise to them from time $t=0$ to $t=1$, arriving at a distribution that is easy to approximate $p(X_{1}).$ Then we train the de-noiser to generate paths that match the paths of the forward process by optimizing an evidence lower bound (ELBO) as
$\log q_{\theta}(X_{0})\geq\mathbb{E}_{p((X_{t})_{t=0}^{1}|X_{0})}\log\frac{q_{\theta}((X_{t})_{t=0}^{1})}{p((X_{t})_{t=0}^{1}|X_{0})}.$ (1)
Typically, however, the forward noising process is chosen to be random substitutions. Accordingly, the de-noiser $q_{\theta}$ only applies substitutions rather than deletions. To search over the space of deletions, we therefore need a new diffusion framework.
3 Related work
Diffusion models with insertions and deletions   In chemistry and language modeling, there have been diffusion models that have attempted to allow for insertions and deletions. Campbell2023-sv propose TDDM, a jump diffusion model to handle varying dimensionality, and Patel2025-kn train a TDDM model on language. Their forward noising process involves randomly deleting elements, such that the stationary distribution is an empty sequence. This allows them to train a model which can learn to expand sequences. As well, Johnson2021-gq formulate a discrete-time noising process for small-scale language modeling that includes insertions, deletions, and substitutions. Unfortunately their loss computation scales with the number of discrete time-steps and it is unclear how to extend their framework to continuous-time diffusion, which is known to be superior [campbell2022continuous]. Contemporary with this work, like Johnson2021-gq, Havasi2025-yx use auxiliary tokens to describe a flow-matching method for training a model to perform insertions and deletions in language.
Compared to these models, we train a diffusion model with inductive biases for the shrinking task. We also build models with competitive likelihoods to other generative models; in particular, unlike the models above (except Johnson2021-gq) our model has a closed-form ELBO, which allows principled model comparisons. To enable model comparison, we solve several mathematical and practical problems:
  • We prove that a formal stationary distribution is not necessary to define a diffusion model (Thm. 4.1). This allows us to train a diffusion model that only learns to shrink.
  • We extend the derivation in Amin2024-xu to derive a “schedule conditioned” loss. This is more stable than classical discrete diffusion, and also allows us to condition on the number of deletions $M$ (Thm. 4.2).
  • We Rao-Blackwellize our gradient estimator by integrating over all insertion paths. We do so in practice by noting a connection with sequence alignments (Prop. 4.3).
  • We solve a number of engineering challenges to 1) pick an optimal rate function, 2) learn on very long sequences, and 3) leverage pretrained weights from ESM (Sec. 4.3).
Leveraging evolutionary data to shrink proteins   Recently, Raygun [Devkota2024-za] also suggested using a model trained on sequences from nature to shrink proteins. Raygun trains a stochastic autoencoder to embed and generate sequences of any length on the UniRef dataset which they apply to a variety of downstream tasks, including shrinking long proteins by decoding their embeddings at a shorter length. However, Raygun cannot enforce similarity between the shrunken sequence and original sequence. Furthermore, like previous generative models of protein sequences, Raygun was not specifically trained to shrink. Below we show that our model, SCISOR, suggests shrunken proteins that more often preserve structure and other indicators of function than Raygun.
4 A diffusion model that learns to delete: SCISOR
We now introduce SCISOR, a diffusion approach that learns to effectively shrink even huge sequences, while preserving function. To search the space of deletions and train a model with the right inductive biases, in Sec. 4.1 we build a process which noises sequences by adding random insertions. Then in Sec. 4.2 we show how to train a de-noiser $q_{\theta}$ that reverses this process (Fig. 1a). Finally, in Sec. 4.3, we discuss the practical choices we made to efficiently train SCISOR. In Sec. 5 we describe how to use the de-noiser to generate sequences, shrink proteins, and plan deletions in practice.
4.1 Forward noising with the pure birth process
We propose an insertion-only forward noising process for discrete diffusion known as the pure birth process [Kendall1948-uu] with rate function $\beta(t)$ and insertion distribution $\pi$. Let $X_{0}$ be a sequence $X_{0}^{(1)},\ldots,X_{0}^{(L)}$. There are $L+1$ possible locations we can insert letters. In the pure birth process, at instant $t$, each of these locations gains an insertion with rate $\beta(t)$. The letter that is inserted is drawn from some distribution $Y\sim\mathrm{Cat}(\pi).$ After $Y$ is inserted at some position, the process continues and there are now $L+2$ positions in which there could be insertions with rate $\beta(t)$ (Fig. 1a). To train a diffusion model to reverse this process, we need to (1) easily sample $p(X_{t}|X_{0})$ and (2) easily approximate $p(X_{1}|X_{0})$.
Sampling $X_{t}$   Rather than simulate the pure birth process up until time $t$, we show in App. F.1 that $X_{t}$ can be sampled directly from $X_{0}$ as in Alg. 1. Note that $0<\alpha(t)\leq 1$ controls how many insertions are added: by the property of negative binomial distributions, the expected length of $X_{t}$ is $\mathbb{E}[M_{t}+L]=\frac{L+1}{\alpha(t)}-1$, which grows as $\alpha(t)\rightarrow 0$.
Algorithm 1 Sample $X_{t}$
0: Initial sequence $X=X^{(1)}\cdots X^{(L)}$, time $t$
1: Compute the probability of no insertions at a site $\alpha(t)\leftarrow\exp\left(-\int_{0}^{t}\beta(s)\,ds\right)$
2: Sample total number of insertions up to time $t$, $M_{t}\sim\text{NegativeBinomial}(L+1,\alpha(t))$
3: Sample the number of insertions in each position by uniformly distributing $M_{t}$ into $L+1$ bins: $(\ell_{0},\dots,\ell_{L})\sim\text{UniformMultinomial}(M_{t})$
4:for $j=0$ to $L$ do
5:  Sample insertion $Y_{j}$ of length $\ell_{j}$, with each character independently from $\text{Cat}(\pi)$
6: Add insertions into $X$ to construct $X_{t}\leftarrow Y_{0}X^{(0)}Y_{1}X^{(1)}\cdots X^{(L)}Y_{L}$
7:return $X_{t}$
Approximating $p(X_{1}|X_{0})$   As $t$ grows, $X_{t}$ becomes longer. To build a diffusion model however, the distribution $p(X_{t}|X_{0})$ typically must converge to a distribution so that it can be approximated by a distribution that can easily be sampled from, $q(X_{1})$. Our critical insight is that $p(X_{t}|X_{0})$, while not converging, can still be very well approximated by long random sequences as $t$ gets large.
Theorem 4.1.
(Proof in App. F.2) Say $X_{0}$ is a sequence with length $L$. Call $q(\cdot\mid L)$ a distribution over sequences of length $L$ which simply samples each letter independently from $\mathrm{Cat}(\pi).$ Then, as the number of insertions increases, $M_{1}\to\infty$, $X_{1}$ becomes easier to approximate with $q$:
$\mathrm{KL}(p(X_{1}\mid X_{0},M_{1})||q(X_{1}\mid L+M_{1}))\to 0.$ (2)
4.2 Learning to reverse this insertion-only noising process
Given a forward process of insertions, we now wish to learn a de-noiser $q_{\theta}$ that generates sequences that resemble those found in nature by deleting letters from long random sequences. We now (1) describe our reverse process $q_{\theta}((X_{t})_{t=1}^{0})$, (2) write the ELBO in Eqn. 1 for our model, and (3) describe how the de-noiser $q_{\theta}$ is being trained toward a target that deletes letters that are unlikely to align with the starting sequence $X_{0}$.
The reverse process   For a forward path $(X_{t})_{t=0}^{1}$ from a sequence $X_{0}$ of length $L$, define $t_{1},\dots,t_{M_{1}}$ to be the times of each insertion. We can then sample forward paths by first deciding how many insertions will occur until time $1$ and when these insertions will occur, and then choosing what these insertions are:
$p((X_{t})_{t=0}^{1}|X_{0})=p(M_{1}|L)p(t_{1},\dots,t_{M_{1}}\mid M_{1},L)\prod_{M=1}^{M_{1}}p(X_{t_{M}}|X_{t_{M-1}}).$
We follow the discrete diffusion framework in Amin2024-xu in defining the reverse process to match the noise schedule of the forward process. To generate a sequence of length $L$, we first decide the number of insertions and their times from the same distribution as $p$, and then denoise each insertion [^2] Note our process is conditioned on generating a sequence of a particular length $L$. ,
$q_{\theta}((X_{t})_{t=1}^{0}|L)=p(M_{1}|L)p(t_{1},\dots,t_{M_{1}}\mid M_{1},L)q(X_{1}|L+M_{1})\prod_{M=1}^{M_{1}}q_{\theta}(X_{t_{M-1}}|X_{t_{M}},M).$
Now we must only train our de-noiser $q_{\theta}(X_{t_{M-1}}|X_{t_{M}},M)$ to take in a sequence $X_{t_{M}}$ and the number of insertions that sequence has $M$, and predict the sequence before the last insertion $X_{t_{M-1}}$. That is, $q_{\theta}(\cdot\mid X_{t_{M}},M)$ can be thought of as a distribution over the letters of $X_{t}$.
The loss   To train the de-noiser, we modify the calculation of the ELBO from Eqn. 1 as in Amin2024-xu. We will then use this ELBO as our objective for training the de-noiser.
Theorem 4.2.
(Proof in App. F.3) Define $M_{t}$ as the number of mutations up to time $t$, and $\mathrm{prev}(X_{t})$ is the last sequence that gained an insertion to become $X_{t}$. Then the negative log likelihood of a sequence of length $L$, $\log q_{\theta}(X_{0}|L)$, is smaller than
$\mathbb{E}_{M_{1}}\mathrm{KL}$ $(p(X_{1}\mid X_{0},M_{1})||q(X_{1}|L+M_{1}))$ (3)
$+\mathbb{E}_{t,X_{t},M_{t}}\frac{M_{t}\beta(t)}{1-\alpha(t)}\mathrm{KL}(p(\mathrm{prev}(X_{t})\mid X_{0},X_{t},M_{t})||q_{\theta}(\mathrm{prev}(X_{t})\mid X_{t},M_{t}))$
The first term is the quantity in Eqn. 2 – how well we can approximate $p(X_{1});$ it is small as long as $M_{1}$ is typically large, i.e. $\alpha(1)$ is small, and can be calculated as in App. A. The second term is the quantity we use to train the de-noiser. $q_{\theta}$ takes in $X_{t}$ and the number of insertions in $X_{t}$ and must predict which letter of $X_{t}$ was last inserted – $\mathrm{prev}(X_{t})$. To train the model, we must calculate $p(\mathrm{prev}(X_{t})|X_{0},X_{t},M_{t})$.
Figure 2: To calculate our target distribution of what letter to delete, $p(\mathrm{prev}(X_{t})\mid X_{0},X_{t},M_{t})$, we align our starting sequence $X_{0}$ to our noised sequence $X_{t}$. The reverse process should favor deleting letters that are gaps in more of the alignments.
Target distribution   Eqn. 3 trains $q_{\theta}$ to match $p(\mathrm{prev}(X_{t})|X_{0},X_{t},M_{t})$, the true distribution over which letter of $X_{t}$ was last inserted in the forward process.
Conditioned on $X_{0},X_{t},M_{t}$, we could find $\mathrm{prev}(X_{t})$ by simulating a pure birth process path from $X_{0}$ to $X_{t}$ and seeing what insertion occurred last. However there are multiple paths that could lead from $X_{0}$ to $X_{t}$; to calculate $p(\mathrm{prev}(X_{t})\mid X_{0},X_{t},M_{t})$, we must marginalize over all of these paths.
The next proposition shows that we can integrate over all of these paths by first enumerating every way to align $X_{0}$ to $X_{t}$ and noting that letters that align with $X_{0}$ less often are more likely to have been $\mathrm{prev}(X_{t})$ (Fig. 2).
Proposition 4.3.
(Proof in App. F.4) Call $\mathrm{ali}(X,Y)$ the number of ways to align a sequence $X$ to a sequence $Y$. Call $b$ the letter that was deleted from $X_{t}$ to $\mathrm{prev}(X_{t})$.
$p(\mathrm{prev}(X_{t})|X_{0},X_{t},M_{t})=\frac{\mathrm{ali}(X_{0},\mathrm{prev}(X_{t}))}{M_{t}\cdot\mathrm{ali}(X_{0},X_{t})}.$
Naively computing this quantity would require running an expensive alignment for every deletion. In practice, we use dynamic programming to computes all $\mathrm{ali}(X_{0},\mathrm{prev}(X_{t}))$ in parallel (App. G).
4.3 Implementing SCISOR at UniRef scale
Our ultimate goal is to train large SCISOR models on huge protein data – in particular, the UniRef database [uniref]. We train the SCISOR de-noiser with mini-batch gradient descent on the second term of Eqn. 3 with i.i.d. samples of $t\sim\mathrm{Uniform}(0,1),X_{0},M_{t},X_{t}$. We now discuss how we choose the rate function $\beta(t)$, the distribution of insertion letters $\pi$, the architecture for $q_{\theta}$, and methods to handle the large variation in sequence lengths of $X_{t}$, which we must pass to $q_{\theta}$.
Hyperparameters   Our choice of hyperparameters follows that of standard diffusion methods. As in Austin2021-dg, Amin2024-xu, we chose the rate function $\beta(t)$ so that the mutual information between $X_{t}$ and $X_{0}$ decreases roughly linearly on the interval $t\in[0,1]$. We then modulated $\beta$ so that $\alpha(1)$ was large enough that the first term of Eqn. 3 is small, while samples in the second term did not get to many very long $X_{t}$. We provide details in App. A. We chose the categorical distribution $\pi$ to match the prevalence of amino acids in our training set.
Architecture   We chose our architecture of the de-noiser $q_{\theta}(\cdot|X_{t},M)$ to leverage the pre-trained weights of a BERT-style protein language model, while modifying the architecture to also condition on $M$. The ESM2 architectures [Lin2022-by] are trained on a masked language modeling task, taking in sequences and outputting logits at every site. We finetuned these models for $q_{\theta}$ by replacing their last layer with a linear and softmax layer. To condition on $M$, we add FiLM layers [Perez2017-tg] between attention blocks, such that coordinate $d$ of the activations in layer $\ell$, $a^{\ell}_{d}$, is modified with an affine linear transformation $(1+A_{\theta,d}^{\ell}(M))\times a^{\ell}_{d}+B_{\theta,d}^{\ell}(M)$, where $A_{\theta}$ and $B_{\theta}$ are shallow fully-connected networks initialized to zero.
Engineering for long sequences   Since $X_{t}$ sequences can have wildly different lengths, training naively could result in passing batches with a high proportion of padding and passing very long sequences into the model. To reduce the proportion of compute spent on padding while maintaining an unbiased estimate of the loss, we sort the $X_{t}$ sequences within a given batch by length, and pass them into the model in smaller sub-batches with accumulated gradients. Next, to handle cases with extremely long $X_{t}$, if $|X_{t}|>2048$, we randomly select a window $X^{(w:w+2048)}_{t}$ uniformly at random to pass to the model. We then re-normalize the model predictions by $2048/|X_{t}|$ and use uniform predictions outside the window such that the deletion probabilities sum to 1. This choice keeps our ELBO a valid lower bound on the likelihood. We provide further details on the ELBO and sampling in App. A.
5 Using SCISOR to generate and shrink protein sequences
The SCISOR de-noiser $q_{\theta}$ is trained as a generative model of natural sequences. In this section, we describe how to unconditionally generate natural sequences. Then we describe the statistical basis by which we may use the de-noiser for downstream tasks: predicting the effect of deletions on a protein’s function, and, shrink long sequences to produce shorter natural sequences.
Figure 3: The SCISOR de-noiser $q_{\theta}$ plans deletions to arrive at sequences that resemble those in nature, and therefore avoids deleting important structural motifs in natural sequences. (a) SCISOR unconditionally samples proteins by starting with a large random sequence $X_{1}$ (light blue) and iteratively deleting according to $q_{\theta}(\mathrm{prev}(X)|X,M)$, to arrive at a protein that resembles those in nature (dark blue). We predict the structure of each sequence with OmegaFold [Wu2022-ma]. (b) We ask SCISOR to plan the first of $M$ mutations for R4SNK4 and color residue $i$ on a structure from Aleku2016-jw by the deletion probability $q_{\theta}(X^{(-i)}|X,M)$. As $M$ increases, SCISOR favors deletions (red) in more regions while minimizing deletions in the catalytic structural motif near the bottom (white).
High-quality unconditional generation   As described in Sec. 4.2, to sample a sequence of length $L$ from SCISOR, one samples a long random sequence from $\mathbb{E}_{M_{1}\mid L}q(X_{1}|L+M_{1})$ and then iteratively deletes according to the de-noiser $q_{\theta}$ (Fig. 3a). Campbell2023-sv suggests continuous-time discrete diffusion models can get higher quality samples, sacrificing some compute, by applying corrector steps which noise and de-noise repeatedly. For SCISOR, this process takes the form of adding and removing insertions as in Alg. 2. This allows SCISOR to more thoroughly search the space of deletions, potentially escaping local minima. In cases where many passes through the model is too expensive, we can make multiple deletions per de-noiser prediction, as we discuss in App. A.
Algorithm 2 Unconditional sequence generation with corrector steps
0: Desired sequence length $L$, corrector steps $K$
1: Sample $M\sim\text{NegativeBinomial}(L+1,\,\alpha(1))$
2: Sample $X$ of length $L+M$ where each $X^{(j)}\sim\text{Cat}(\pi)$ independently
3:while $|X|>L$ do
4:  for $k=1$ to $K$ do
5:   Remove one letter from $X$ according to $q_{\theta}(\mathrm{prev}(X)\mid X,M)$
6:   Insert a random letter from the distribution $\pi$ into a random position in $X$
7:  Remove one letter from $X$ according to $q_{\theta}(\mathrm{prev}(X)\mid X,M)$
8:  $M\leftarrow M-1$
9:return $X$
Mutation effect prediction   Say we have a sequence $X$ and we wish to predict the effect of the deletion of every position to understand the importance of each residue. Typically, we would take a model trained on protein sequences, $p_{\theta}$ and then evaluate the naturalness of the sequence with each deletion through the likelihood $p_{\theta}(X^{(-i)})$, where $X^{(-i)}$ is the deletion of letter $i$ [Riesselman2018-nm]. Unfortunately estimating the likelihood is challenging for diffusion models as one needs to estimate the expectation in Eqn. 1.
SCISOR instead simply predicts $q_{\theta}(X^{(-i)}\mid X,M=1)$ for every possible deletion $X^{(-i)}$. Then if the de-noiser suggests that a residue is unlikely to be deleted, then $X$ without that residue does not look like a sample from $q_{\theta}(X_{0})$, i.e. a natural sequence, and thus that deletion may harm function. For multi-letter deletions, we integrate over all deletion paths (see App. A).
[…]
Note that SCISOR is not simply sampling deletions by how natural $\mathrm{prev}(X)$ looks. Rather it also uses knowledge of $M$ to plan for future mutations. Different values of $M$ allow the model to change which deletions it will allow at each step (Fig. 3b).
Given enough data and compute, SCISOR should learn the correct distribution, $q_{\theta}(X_{0}=\tilde{X}\mid X,M)$. However, in principle, there is a distribution shift between sampling $q_{\theta}(X_{0}=\tilde{X}\mid X,M)$ when $X$ is a realistic protein versus when $X$ is a sequence with noisy insertions in our learning process, which may impact the statistical efficiency of the learning process. In practice, App. C and our results in the following sections show that the SCISOR de-noiser learns meaningful evolutionary signals.
6 SCISOR is a competitive generative model for proteins
Figure 4: SCISOR fits the distribution of sequences in nature competitively with established sequence modeling approaches. (a) SCISOR is competitive with other diffusion models (grey) in perplexity. "S, M, L" refer to model size. (b, c) Samples from SCISOR ($K=5$) are predicted to be competitive quality to those from diffusion models and competitive with AR models as measured by (b) matching the distribution of natural sequences as measured by the Fréchet protein distance (FPD) and (c) foldability (higher pLDDT from OmegaFold [Wu2022-ma]). We took EvoDiff and AR perplexities from Alamdari2023-nj.
We now evaluate how well SCISOR fits the distribution of natural sequences compared to established sequence modeling methods. All details of our experimental setup are in App. B.
In Fig. 4, we compare the quality of SCISOR’s fit to the data against state-of-the-art protein diffusion models: EvoDiff [Alamdari2023-nj] and DPLM [Wang2024-bn]. It is well known that diffusion models regularly under-perform autoregressive models on fitting the data; we therefore include two autoregressive models from Alamdari2023-nj as references. All models are trained on the same release of UniRef50 [uniref] – small models have 35-38M parameters, DPLM M has 150M parameters, and large models have 640-650M parameters. We evaluate each model’s perplexity on a test set, and the quality of their samples, as measured by how well they match the distribution of natural sequences (FPD), and foldability (pLDDT).
Despite its difference from established modeling methods, SCISOR is competitive with other diffusion models in perplexities. Moreover, SCISOR often generates higher-quality samples than previous diffusion models, even competitive with the AR reference, likely due to SCISOR being a continuous-time model, as we discuss in Sec. 5.
7 State-of-the-art functional effect predictions for deletions
We evaluate SCISOR’s ability to predict the effect of deletion mutations on the function of proteins as measured in the lab. To do so, we use 7000 measurements of deletion effects from 62 assays collected in ProteinGym [Notin2023-eh]. As baselines, we compare against existing mutation effect prediction models, including state-of-the-art autoregressive models ProGen2 [Nijkamp2022-ip] and Tranception [Notin2022-qw]. Since models trained on UniRef90 tend to better predict the effects of mutations [Rives2021-co], all models in this section are trained on UniRef90.
In Fig. 5, we report the Spearman correlations of the measurements of each assay against the predicted effects from each model, taking the best performance within each model family (full table in App. E). SCISOR outperforms all baselines on both the single-deletion and multi-deletion benchmarks, even outperforming PoET [Truong2023-dk], a large model that has access to extra information about protein families.
Figure 5: SCISOR makes state-of-the-art predictions for the effect of deletions on protein function measured in the lab. We calculate the average Spearman correlation between predicted deletion effects and measurements across all assays in ProteinGym, presenting the results from the highest-performing variant of each model architecture. Models that use multiple sequence alignment information are striped.
8 Preserving key in-silico indicators of function when shrinking
Figure 6: SCISOR shrinks proteins while preserving their key properties. We take 200 UniProt sequences with binding or active site annotations and shrink them by various amounts. We measure (a) pLDDT structural confidence scores from Omegafold, predicted structural similarity between the original and shrunk proteins with (b) TM and (c) RMSD, and (d) conservation of annotated functional sites, measured by the enrichment ratio of annotated sites relative to random expectation. Horizontal dashed lines indicate reference values for 0% shrinking.
We now evaluate the ability of the SCISOR de-noiser $q_{\theta}$ to propose promising shrunk samples of long proteins from nature. Specifically, we take 200 sequences from UniProt that have binding or active site annotations and shrink them to various amounts.
Since the diffusion models in Sec. 6 cannot suggest deletions, we compare as baselines to shrinking with ProGen2 and Raygun. Raygun requires $1$ model evaluation to make $M$ deletions, while SCISOR requires $M$. For ProGen2, ideally we would sample substrings of $X$ with length LML-M proportional to ProGen2’s likelihood; however, that approach would require $\binom{L}{M}$ model evaluations, which is prohibitively expensive. We instead consider a strong but tractable baseline: we predict the effect of all $L$ single deletions independently (in $L$ model evaluations) and then sample sets of deletions with probabilities based on these effects. Further details are in appendix˜B and section˜E.3.
In Fig. 6 we see SCISOR consistently suggests shrunken proteins that are more likely to be foldable and preserve key indicators of function (predicted structural topology, presence of active or binding sites) than Raygun and ProGen2 baselines. Although ProGen2 achieves higher functional site preservation when shrinking by 50%, the lower pLDDT scores suggest that its samples at this level of shrinkage are not likely to be functional. In App. E.2, we show that SCISOR also achieves best performance for greedy sampling of shrunk sequences, reflecting situations in which a practitioner is not interested in generating diverse samples. Lastly, in appendix˜D we perform an in-depth case study of shrinking the GTP sensor RalA, showing that shrinking with SCISOR best preserves the predicted structure of the binding site with GTP.
9 Conclusion
By proposing a new family of generative models that learn to build natural sequences by deleting, SCISOR, we have built models that can effectively shrink proteins. Future work may seek to address some of the conceptual limitations of the SCISOR process.
Realistic insertion process   One way to mitigate the distribution shift mentioned in Sec. 5 is to make samples from $p(X_{t})$ look more like natural sequences with a more elaborate forward process. The challenge is finding a process that provably achieves an easy-to-approximate $p(X_{1})$ as in Sec. 4.1 and deriving a closed-form integral over all paths as in Sec. 4.2. Future work may leverage our theoretical and practical advances to derive the losses and train such models at large scale.
Guiding based on function   In this work, we aimed to shrink proteins into sequences that may still appear in nature and are thus likely to be functional. While two functional proteins with similar sequences are likely to have the same function, this is not guaranteed, especially in those protein families with diverse functions [Zhang2024-tb]. Future work may incorporate other information of function into the SCISOR shrinking process. For example, one could guide the SCISOR diffusion process using a classifier trained to detect functional proteins of interest [Nisonoff2024-bs].
Including compensatory mutations   Currently, SCISOR only shrinks proteins via deletions. It is possible however that there are substitutions or insertions that could be added to a protein to make it more tolerant to more deletions. To allow SCISOR to introduce these mutations when planning a series of deletions, we could add substitutions and deletions to the forward process, and train the de-noiser to also include substitutions and insertions in its planning.
Acknowledgements
We thank Alex Ali and Andres Potapczynski for helpful feedback. This work was supported in part by NSF CAREER IIS-2145492, NSF CDS&E- MSS 2134216, NSF HDR-2118310, BigHat Biosciences, and Capital One.
Appendix A Details about SCISOR
A.1 Prior matching KL term
We rewrite the first term of Eqn. 3 so we can estimate it.
Proposition A.1.
(Proof in App. F) $\mathrm{KL}(p(X_{1}\mid X_{0},M_{1})||q(X_{1}|L+M_{1}))$ is equal to
$\mathbb{E}_{X_{1}\mid X_{0},M_{1}}\left[\log\binom{M_{1}+L}{L}+\sum_{i=1}^{L}\log\pi(X_{0}^{(i)})-\log\mathrm{ali}(X_{0},X_{1})\right].$ (4)
We can therefore estimate the first term of the loss in Eqn. 3 by sampling $X_{0},M_{1},X_{t}$ as calculating the quantity in the expectation of Eqn. 4.
A.2 Efficient sampling
Alg. 2 implements the Gillespie algorithm for a stochastic process. Zhao2024-fv and Amin2024-xu suggested $k$-Gillespie for diffusion models, taking $k$ steps at every step by sampling without replacement. Indeed we can do the same for SCISOR, sampling many deletions at each step without replacement.
A.3 Multi-deletion prediction
Say $\tilde{X}$ is the sequence $X$ with $M$ deletions at sites $\{i_{1},\dots,i_{M}\}.$ We wish to calculate $q_{\theta}(X_{0}=\tilde{X}\mid X,M).$ We can break this up into a sum over all deletions using the de-noiser
$q_{\theta}(X_{0}=\tilde{X}\mid X,M)=\sum_{m=1}^{M}q_{\theta}(X_{0}=\tilde{X}\mid X^{(-i_{m})},M-1)q_{\theta}(\mathrm{prev}(X)=X^{(-i_{m})}\mid X,M).$
Continuing like this, we can write $q_{\theta}(X_{0}=\tilde{X}\mid X,M)$ as a sum over all permutations of the deletions.
Algorithm 3 Predicting the functional effect of multiple deletions with SCISOR
0: Initial sequence $X$, deletions $\{i_{1},\dots,i_{M}\}$.
[…]
3:for $j_{1},\dots,j_{M}\in P$ do
4:  $\mathrm{SUM}=\mathrm{SUM}+\prod_{M^{\prime}=0}^{M-1}q_{\theta}(X^{(-j_{1},\dots,j_{M^{\prime}+1})}\mid X^{(-j_{1},\dots,j_{M^{\prime}})},M^{\prime})$
5:return $\mathrm{SUM}=q_{\theta}(X_{0}=\tilde{X}\mid X,M)$
A.4 Rate function
For simplicity, we choose a functional form
$\beta(t)=\frac{\gamma}{1-t_{\mathrm{max}}t}.$
Consequently, we have:
$\alpha(t)=$ $\mathrm{exp}\left(-\int_{0}^{t}\beta(s)\,ds\right)$
$=$ $\mathrm{exp}\left(-\frac{\gamma}{t_{\mathrm{max}}}\int_{0}^{t}\frac{1}{1-t_{\mathrm{max}}s}\,ds\right)$
$=$ $\mathrm{exp}\left(-\frac{\gamma}{t_{\mathrm{max}}}\ln(1-t_{\mathrm{max}}t)\right)$
$=$ $(1-t_{\mathrm{max}}t)^{\gamma/t_{\mathrm{max}}}$
and $\alpha(1)=(1-t_{\mathrm{max}})^{\gamma/t_{\mathrm{max}}}.$ Now we must choose $\gamma$ and $t_{\mathrm{max}}$. We found empirically on small models that $\gamma=1.1$ gave an ELBO (see Eq.  3) such that the expectation conditional on each $t$ was roughly even. We found empirically on small models that $t_{\mathrm{max}}=0.9$ gave the best best loss controlling for wall time, trading off allowing the model to attempt to fit larger sequences and spending compute on those large sequences.
A.5 Windowing
One challenge in efficiently training the SCISOR de-noiser is that we must compute $q_{\theta}(\mathrm{prev}(X_{t})\mid X_{t},M_{t})$, where $X_{t}$ can potentially be a very long sequence. To handle these long sequences, we introduce a windowing strategy: if $|X_{t}|>2048$, we randomly select a window $X^{(w:w+2048)}_{t}$ uniformly at random to pass to the model. We then re-normalize the model predictions by $2048/|X_{t}|$ (the probability of a deletion in the window is proportional to its size) and use uniform predictions outside the window such that the deletion probabilities sum to 1. Calling the predictions made by window $w$ $q_{\theta}^{w}(\mathrm{prev}(X_{t})\mid X_{t},M_{t})$, we can define our model predictions as an average over all windows
$q_{\theta}(\mathrm{prev}(X_{t})\mid X_{t},M_{t})=\mathbb{E}_{w}q_{\theta}^{w}(\mathrm{prev}(X_{t})\mid X_{t},M_{t}).$
ELBO
We modify the second term of our loss Eqn. 3 to obtain another lower bound to bring the expectation outside
$\mathrm{KL}(p(\mathrm{prev}(X_{t})\mid X_{0},X_{t},M_{t}$ $)||q_{\theta}(\mathrm{prev}(X_{t})\mid X_{t},M_{t}))$
$\geq$ $\mathbb{E}_{w}\mathrm{KL}(p(\mathrm{prev}(X_{t})\mid X_{0},X_{t},M_{t})||q^{w}_{\theta}(\mathrm{prev}(X_{t})\mid X_{t},M_{t})).$
This gives us a new ELBO we can estimate by stochastically sampling the window $w$ whenever we get a large sequence.
Sampling
In Alg. 1, we need to sample from $q_{\theta}(\mathrm{prev}(X_{t})\mid X_{t},M_{t})$ for very long sequences. We do so by sampling a $w$ and then sampling from $q_{\theta}^{w}(\mathrm{prev}(X_{t})\mid X_{t},M_{t})$.
Appendix B Experimental Details
B.1 Baselines
We used EvoDiff models and code from https://github.com/microsoft/evodiff under the MIT license. We used DPLM models and code from https://github.com/bytedance/dplm under the Apache-2.0 license. We used ProGen2 models and code from https://github.com/enijkamp/progen2 under the BSD-3-clause license. We used Raygun models and code from https://github.com/rohitsinghlab/raygun under the CC BY-NC 4.0 license. We used ProteinGym models and code from https://github.com/OATML-Markslab/ProteinGym under the MIT license.
B.2 SCISOR architecture
We used the flash attention implementation of ESM from faesm2024 under the MIT license. We used ESM2 weights [Lin2022-by] also under the MIT license. We developed SCISOR using code from https://github.com/AlanNawzadAmin/SCUD under the MIT license.
B.3 Training SCISOR
We apply our framework to train a protein generative model on UniRef50 [uniref]. We filter this dataset to exclude proteins with non-standard amino acids, and crop long protein sequences down to their first 1024 amino acids.
For the results in section˜6, we train SCISOR models on the March 2020 release of UniRef50, using the same train-test split as EvoDiff [Alamdari2023-nj] from https://zenodo.org/records/6564798. Our models were trained about one week each on one NVIDIA A100 GPU with an effective batch size of 256 and learning rate of 0.0001.
For the results in sections˜8 and 7, we train SCISOR models on the latest release of UniRef90. Here, we use an effective batch size of 512 and learning rate of 0.00005. The SCISOR S and M models were trained for about one week each on two NVIDIA A100 GPUs. The SCISOR L model was trained for about four days on four NVIDIA H100 GPUs.
For each effective batch, we sampled all $t,X_{0},M_{t},X_{t}$. We then sorted sequences by the length of $X_{t}$ before breaking them into batches to pass to the model in batch sizes of $8$ or $16$; This makes sequences in each batch have similar length, minimizing padding.
B.4 Model fit experiments
B.4.1 Perplexities
SCISOR
We compute the perplexity in Fig.  4a on the test dataset by first sub-sampling the expectation of the ELBO from Prop. 4.2 – we take 10 samples of $t,X_{t}$ for every sequence. We then by the total number of tokens in the test set and report the exponentiated negative result.
EvoDiff and AR
We take perplexity values from Table S1 in  Alamdari2023-nj.
DPLM
DPLM was trained as a discrete-time masking diffusion model with $500$ steps and a linear rate schedule – that is, the probability of each token in $X_{t}$ being masked is $t/500$. We therefore evaluated their perplexities as such a model as in Austin2021-dg. This ELBO becomes
$\sum_{t=1}^{500}\frac{1}{t}\mathbb{E}_{X_{0},X_{t}}\sum_{i=1}^{L}\mathbbm{1}(X_{t}^{(i)}=\mathrm{mask})\log q_{\theta}(X_{0}^{(i)}\mid X_{t}).$
B.4.2 Samples
SCISOR
We sampled according to Alg. 2.
EvoDiff and AR
We sampled from EvoDiff and AR models using functions generate_oaardm and generate_autoreg from https://github.com/microsoft/evodiff/blob/main/evodiff/generate.py.
DPLM
Wang2024-bn suggested a novel sampling method for DPLM. However, we were interested in measuring the quality of DPLM samples as a diffusion model. We therefore took samples as such a model as in Austin2021-dg: We start with $X_{500}$ and for every $t=500,\cdots,1$ we unmask each position $i$ with probability $1/t$, replacing the mask according to predicted probabilities $q_{\theta}(X_{0}^{(i)}\mid X_{t})$.
B.4.3 Sample evaluation
For FPD we took 1000 protein lengths from UniRef50 and sampled sequences of each of those lengths from SCISOR, EvoDiff, and DPLM; or we sampled 1000 sequences from the AR models. For pLDDT, we sampled 100 sequences of length 100, 200, 400, and 800 from SCISOR, EvoDiff, and DPLM; for AR models where the sample length cannot be controlled, we sampled sequences until we had a sufficient number of samples with lengths within $10\%$ of each desired length.
Fréchet protein distance (FPD)
We calculated the FPD of 1000 generated sequences to 10000 samples from UniRef50 using ProtT5 embeddings in https://github.com/hefeda/PGP under the Apache-2.0 license. We then calculated the Fréchet inception distance between the embeddings of the natural sequences and each set of sampled sequences as
$\|\mu_{\mathrm{natural}}-\mu_{\mathrm{sample}}\|^{2}+\mathrm{tr}\left(\Sigma_{\mathrm{natural}}+\Sigma_{\mathrm{sample}}-2(\Sigma_{\mathrm{natural}}\Sigma_{\mathrm{sample}})^{1/2}\right)$
where $\mu_{\cdot}$ and $\Sigma_{\cdot}$ are empirical means and covariances of the embeddings.
pLDDT
We calculate pLDDT scores using OmegaFold [Wu2022-ma] as described in https://github.com/HeliXonProtein/OmegaFold/blob/main/README.md under the Apache-2.0 License. For computational efficiency, we use only 1 cycle per sample. This results in lower overall pLDDT scores than the recommended default settings, which uses 10 cycles to obtain more accurate predicted structures.
B.5 ProteinGym
B.5.1 Model predictions
SCISOR
To evaluate SCISOR, we set $X_{t}$ to be the target sequence and $M$ to be the number of deletions between the target and the mutant of interest. We then predict the effect of the deletion using Alg. 3.
ProGen and other models
We evaluated other models using scripts available on ProteinGym.
B.5.2 Model evaluation
For Fig. 5, we adapt the ProteinGym benchmark from [Notin2023-eh] by filtering their indels dataset to cases where the mutant is a strict subsequence of the target sequence. For the single deletions benchmark, we use mutants that are only one deletion away from the target sequence, while for the multiple deletions benchmark, we use mutants that are two or three deletions away from the target sequence.
For single mutations, we gathered $61$ assays in ProteinGym with $4544$ mutations in total.
Three assays in ProteinGym measured double and triple mutations: A4_HUMAN_Seuma_2022 measured stability and had 42 double mutations and 40 triple mutations, KCNJ2_MOUSE_Macdonald_2022 measured expression and had 397 double mutations and 387 triple mutations, P53_HUMAN_Kotler_2018 measured organismal fitness and had 172 double mutations and no triple mutations.
B.6 Shrinking
For Fig. 6a and  6d we sample 100 sequences with annotated active sites and 100 sequences with annotated binding sites from UniProt. We then shrink each sequence by $d$ percent, where $d\in\{1,3,5,10,20,30,40,50\}$.
B.6.1 Model samples
SCISOR
We shrunk sequences using Alg. 2.
ProGen
[…]
$\log q_{\mathrm{ProGen}}(\tilde{X})\approx\log q_{\mathrm{ProGen}}(X)+\sum_{\mathrm{deletions}\ i}\Delta_{i}$
where $\Delta_{i}$ is the effect of a single mutation,
$\Delta_{i}=\log\frac{q_{\mathrm{ProGen}}(X^{(i)})}{q_{\mathrm{ProGen}}(X)}.$
This approximation requires calculating $L$ quantities $\Delta_{i}.$
Sampling from this approximation is equivalent to sampling $M$ deletions – deletion $i$ is sampled with probability proportional to $\exp(\Delta_{i})$. Greedy shrinking just involves picking the $M$ mutations with the highest $\Delta_{i}$.
Raygun
we use the Raygun generate command to generate shrunken proteins of desired length, where length was calculated by first calculating rounded up number of deletions to introduce, and conditioning Raygun to generate sequence of length LML-M. we used a noise ratio of 0.5 with uniform sampling (noise sampled uniformly between 0 and 0.5), in order to limit the number of substitutions introduced. We use a filter ratio of 0.1 meaning we select the best candidate among ten generated sequences, and recycle sequences once.
B.6.2 Model evaluation
We evaluate the foldability of the shrunk sequences using the average pLDDT per residue for the structure generated using OmegaFold [Wu2022-ma] as described in https://github.com/HeliXonProtein/OmegaFold/blob/main/README.md under the Apache-2.0 License, using 1 cycle per sample. We calculate enrichment as the number of active or binding sites in the original sequence that were preserved in the shrunk sequence – we call a functional site “preserved” if no residues were modified or deleted.
Appendix C SCISOR deletions match those in nature for R4SNK4
We collected 5000 sequences locally aligned to R4SNK4 using BLAST [Camacho2009-og] and counted how frequently each position of the R4SNK4 sequence landed in a region that wasn’t aligned (“Not aligned”) or aligned with a gap token (“Gap”). These are two distinct ways to measure deletions made by nature. Next, we shrunk R4SNK4 using SCISOR at various values of $M$, taking 512 samples for each $M$.
In Fig. 7a, we plot, for each position of R4SNK4, how frequently we observe deletions as suggested by SCISOR versus as seen in nature. We see qualitatively that at $M=50$, SCISOR prefers making deletions in regions where deletions are observed in nature, and at $M=100$ SCISOR prefers deleting the second half of the protein, the region which was less frequently included in alignments. This separate behavior at different $M$ is shown quantitatively when we measure the Spearman correlations between SCISOR deletion frequencies and deletions seen in the alignments in Fig. 7b. Almost all comparisons are statistically significant – a random model never achieves a Spearman correlation above 0.015.
Figure 7: SCISOR suggests deletions in R4SNK4 seen in nature. (a) We compare two ways to measure deletions seen in an alignment versus deletions at two values of $M$ from samples of SCISOR. (b) We quantitatively measure the Spearman correlations between the two types of deletions in our alignment to deletions seen in SCISOR samples across $M$.
Appendix D Case Study: Shrinking RalA
To better understand SCISOR’s ability to shrink natural proteins, we perform a detailed case study of the RalA protein, of length 206. We chose RalA since, as a GTP sensor, we could easily tell which sequences were unlikely to be functional by predicting their bound structure to GTP. We evaluate the function of shrunk samples by predicting their binding to GTP as in the structure pdb_00001uad and comparing to RalA. We use Alphafold [alphafold] to predict the structure of shrunken RalA, GTP and binding partner Sec5.
Qualitative results
We measure the TM score (a popular measure of structural similarity to ground truth [Zhang2004-fa]) of wild type and shrunken RalA residues within 10 angstroms of the GTP molecule. Following [xu], we use TM=0.5 as a liberal cutoff for non-function and evaluate 5 designs for each method. The results are presented in Table 1. We see that proteins shrunk with SCISOR are much more likely to adopt the same complex fold as the wild type compared to designs from other methods.
% Deleted Random Raygun ProGen2 SCISOR
1 5 5 5 5
3 4 5 5 5
5 5 5 3 5
10 1 4 1 5
20 0 1 0 5
Table 1: Number of designs not predicted non-functional (TM scores above 0.5) (out of 5)
Quantitative results
Next we show similar results with various measurements of the quality of the interface between the designed RalA and GTP. In Table 2, we perform a structural alignment using the RalA structures and then calculate the RMSD for all residues within 10 angstroms of the center of mass of the GTP. In Table 3, we instead calculate the RMSD of the atoms of the translated and rotated GTP. Additionally, we report DockQ values [dockq] in Table 4 and the ipTM scores from Alphafold in Table 5. In all these results, we we see that SCISOR consistently outperforms all baselines, suggesting that the candidate designs from SCISOR are most likely to preserve the function of the original RalA protein.
% Deleted Random ProGen2 Raygun SCISOR
1 0.90 $\pm$ 0.13 0.87 $\pm$ 0.10 1.02 $\pm$ 0.23 0.92 $\pm$ 0.01
3 1.52 $\pm$ 0.24 1.25 $\pm$ 0.15 0.89 $\pm$ 0.11 0.91 $\pm$ 0.01
5 1.64 $\pm$ 0.15 1.48 $\pm$ 0.37 0.87 $\pm$ 0.17 0.93 $\pm$ 0.00
10 1.83 $\pm$ 0.23 1.93 $\pm$ 0.28 1.03 $\pm$ 0.11 1.24 $\pm$ 0.14
20 2.07 $\pm$ 0.30 2.28 $\pm$ 0.18 2.37 $\pm$ 0.11 1.28 $\pm$ 0.15
Table 2: Interface RMSD of residues within 10 Å of GTP after structural alignment.
% Deleted Random ProGen2 Raygun SCISOR
1 1.69 $\pm$ 0.30 4.03 $\pm$ 1.67 7.39 $\pm$ 4.82 2.07 $\pm$ 0.56
3 3.91 $\pm$ 2.04 6.04 $\pm$ 2.30 3.47 $\pm$ 0.53 2.37 $\pm$ 0.36
5 6.19 $\pm$ 3.20 3.53 $\pm$ 0.18 15.24 $\pm$ 5.37 3.50 $\pm$ 0.72
10 10.92 $\pm$ 5.29 11.47 $\pm$ 6.11 12.19 $\pm$ 4.13 3.95 $\pm$ 1.24
20 17.83 $\pm$ 5.55 10.06 $\pm$ 1.52 15.24 $\pm$ 1.89 10.32 $\pm$ 2.81
Table 3: Ligand RMSD of GTP atoms after structural alignment.
% Deleted Random ProGen2 Raygun SCISOR
1 0.86 $\pm$ 0.04 0.71 $\pm$ 0.17 0.38 $\pm$ 0.15 0.87 $\pm$ 0.01
3 0.50 $\pm$ 0.20 0.51 $\pm$ 0.18 0.62 $\pm$ 0.15 0.85 $\pm$ 0.02
5 0.46 $\pm$ 0.18 0.18 $\pm$ 0.11 0.23 $\pm$ 0.13 0.88 $\pm$ 0.01
10 0.36 $\pm$ 0.15 0.18 $\pm$ 0.15 0.13 $\pm$ 0.10 0.50 $\pm$ 0.18
20 0.13 $\pm$ 0.11 0.11 $\pm$ 0.07 0.03 $\pm$ 0.02 0.30 $\pm$ 0.16
Table 4: DockQ scores for RalA binding to Sec5 with standard error.
% Deleted Random ProGen2 Raygun SCISOR
1 0.83 $\pm$ 0.07 0.82 $\pm$ 0.08 0.60 $\pm$ 0.07 0.90 $\pm$ 0.00
3 0.64 $\pm$ 0.11 0.71 $\pm$ 0.08 0.82 $\pm$ 0.08 0.90 $\pm$ 0.00
5 0.67 $\pm$ 0.09 0.52 $\pm$ 0.08 0.51 $\pm$ 0.01 0.89 $\pm$ 0.00
10 0.40 $\pm$ 0.14 0.47 $\pm$ 0.13 0.59 $\pm$ 0.07 0.67 $\pm$ 0.09
20 0.22 $\pm$ 0.06 0.36 $\pm$ 0.06 0.42 $\pm$ 0.03 0.58 $\pm$ 0.09
Table 5: ipTM scores from AlphaFold for RalA binding Sec5.
Appendix E Supplementary Results
E.1 Full ProteinGym Results
We show the results for ProteinGym for all models and sizes, stratifying the single deletions into functional, taxonomic, and MSA depth categories.
Table 6: ProteinGym results on single and multiple deletions.
Model MSA Single Deletions Multiple Deletions
ProGen2 S 0.457 0.445
ProGen2 M 0.513 0.385
ProGen2 Base 0.497 0.408
ProGen2 L 0.491 0.375
ProGen2 XL 0.393 0.392
RITA S 0.409 0.274
RITA M 0.448 0.318
RITA L 0.465 0.323
RITA XL 0.440 0.161
Tranception S 0.439 0.475
Tranception M 0.464 0.424
Tranception L 0.445 0.426
HMM Yes 0.453 0.474
PoET (200M) Yes 0.551 0.488
SCISOR S 0.332 0.268
SCISOR M 0.505 0.520
SCISOR L 0.573 0.492
Table 7: ProteinGym results on single deletions stratified by the measured function of each assay.
Model MSA Activity Expression Organismal Fitness Stability
ProGen2 S 0.566 0.294 0.499 0.470
ProGen2 M 0.574 0.404 0.558 0.514
ProGen2 Base 0.592 0.380 0.496 0.520
ProGen2 L 0.550 0.344 0.560 0.508
ProGen2 XL 0.418 0.298 0.333 0.521
RITA S 0.507 0.320 0.452 0.356
RITA M 0.514 0.345 0.500 0.432
RITA L 0.530 0.437 0.420 0.474
RITA XL 0.532 0.385 0.360 0.481
Tranception S 0.542 0.351 0.532 0.331
Tranception M 0.594 0.340 0.526 0.395
Tranception L 0.533 0.336 0.445 0.466
HMM Yes 0.496 0.321 0.501 0.493
PoET (200M) Yes 0.664 0.424 0.566 0.551
SCISOR S 0.376 0.289 0.198 0.465
SCISOR M 0.514 0.362 0.576 0.571
SCISOR L 0.604 0.415 0.668 0.606
Table 8: ProteinGym results on single deletions stratified by the MSA depth of proteins in each assay.
Model MSA Low Medium High
ProGen2 S 0.558 0.429 0.497
ProGen2 M 0.415 0.483 0.544
ProGen2 Base 0.438 0.460 0.568
ProGen2 L 0.513 0.473 0.532
ProGen2 XL 0.216 0.499 0.530
RITA S 0.300 0.293 0.424
RITA M 0.278 0.376 0.492
RITA L 0.444 0.434 0.504
RITA XL 0.139 0.462 0.508
Tranception S 0.467 0.316 0.360
Tranception M 0.297 0.358 0.447
Tranception L 0.519 0.391 0.518
HMM Yes 0.624 0.506 0.471
PoET (200M) Yes 0.595 0.553 0.548
SCISOR S 0.385 0.381 0.509
SCISOR M 0.641 0.547 0.575
SCISOR L 0.621 0.628 0.584
Table 9: ProteinGym results on single deletions stratified by the taxa of the protein in each assay.
Model MSA Human Eukaryote Prokaryote Virus
ProGen2 S 0.506 0.467 0.361 0.567
ProGen2 M 0.536 0.539 0.432 0.510
ProGen2 Base 0.568 0.541 0.396 0.471
ProGen2 L 0.536 0.513 0.442 0.492
ProGen2 XL 0.511 0.537 0.420 0.573
RITA S 0.398 0.353 0.272 0.448
RITA M 0.496 0.417 0.310 0.504
RITA L 0.522 0.473 0.327 0.568
RITA XL 0.510 0.466 0.386 0.555
Tranception S 0.332 0.363 0.297 0.449
Tranception M 0.443 0.406 0.295 0.468
Tranception L 0.511 0.462 0.348 0.513
HMM Yes 0.585 0.392 0.437 0.547
PoET (200M) Yes 0.554 0.522 0.523 0.721
SCISOR S 0.486 0.418 0.340 0.652
SCISOR M 0.571 0.539 0.525 0.735
SCISOR L 0.590 0.568 0.621 0.767
E.2 Deterministic Shrinking
In section˜8, we report results for sampling shrunk sequences from SCISOR and baseline models. Those evaluations are most relevant to practical settings, where practitioners may wish to generate a diverse set of promising candidate designs to be evaluated in the lab. Here, we instead consider deterministic shrinking, where the most likely deletions are performed. For Raygun, we mimic this setting by using setting the noise parameter to 0. In fig.˜8, we see that SCISOR consistently achieves excellent performance in this setting as well. Although ProGen2 can achieve higher performance in some cases, this method is more computationally expensive than SCISOR. Moreover, our metrics may be rewarding the “myopic” nature of our ProGen2 baseline, which likely favors deletions that preserve functional regions in the original protein that are not necessarily functional in the shrunken protein.
Figure 8: Deterministically shrinking proteins with SCISOR yields high-quality shrunk samples. All details are as in fig.˜6, except that we take the most likely deletions predicted from each model. For Raygun, we set $\textrm{noise}=0$.
E.3 Computationally Expensive ProGen2 Baseline
In section˜8, we reported a baseline for ProGen2 that required $L$ forward passes of the model, already exceeding the $M$ forward passes used by SCISOR. That baseline assumes that the effects of each deletion are independent, which may be unrealistic. In table˜10, we present results for a more expensive ProGen2 baseline, requiring $\mathcal{O}(L\cdot M)$ forward passes, where deletions are performed one at a time, and the effects of subsequent deletions recalculated. We see that switching to this more expensive ProGen2 baseline does not significantly improve performance.
% Deleted ProGen2 - $\mathcal{O}(L)$ ProGen2 - $\mathcal{O}(LM)$ SCISOR - $\mathcal{O}(M)$
1 71.6 71.7 74.0
3 68.2 67.9 74.2
5 63.5 62.8 74.1
10 55.4 53.0 73.1
20 40.7 36.4 71.2
30 32.3 31.2 63.7
40 31.3 31.9 55.1
50 30.9 33.5 44.7
Table 10: Average pLDDT structural confidence scores from Omegafold for various amounts of shrinkage.
Appendix F Proofs
F.1 Proof of correctness for Algorithm 1
The correctness of Alg. 1 follows from Cor. F.2.
Theorem F.1.
Call
$X_{t}=Y_{0}X_{0}^{(1)}Y_{1}X_{0}^{(2)}\cdots X_{0}^{({L})}Y_{L}$
where $X_{0}^{(1)}X_{0}^{(2)}\cdots X_{0}^{(L)}$ are the letters of $X_{0}$ and $Y_{0},Y_{1},\dots,Y_{L}$ are the insertions. Then $|Y_{l}|$ is a $\mathrm{Geom}(\alpha(t))$ distribution, where $\alpha(t)=\exp(-\int_{0}^{t}\beta(s)ds)$.
Proof.
By the Kolmogorov forward equation,
$\frac{d}{dt}p(|Y_{l}|=n|t)=\beta(t)np(|Y_{l}|=n-1|t)-\beta(t)(n+1)p(|Y_{l}|=n|t).$
This can be written as
$\frac{d}{dt}\left(e^{(n+1)\int_{0}^{t}\beta(s)ds}p(|Y_{l}|=n|t)\right)=ne^{(n+1)\int_{0}^{t}\beta(s)ds}\beta(t)p(|Y_{l}|=n-1|t).$
For $n=0$, this is solved by $p(|Y_{l}|=0|t)=\alpha(t)$. By induction,
$p(|Y_{l}|=n|t)=\alpha(t)(1-\alpha(t))^{n}$
as
$n\alpha(t)^{-(n+1)}\beta(t)p(|Y_{l}|=n-1|t)=$ $n\alpha(t)^{-n}\beta(t)(1-\alpha(t))^{n-1}$ (5)
$=$ $\frac{d}{dt}\left(\alpha(t)^{-n}(1-\alpha(t))^{n}\right)$
Corollary F.2.
$p(|Y_{0}|,\dots,|Y_{L}|)=\alpha(t)^{|X_{0}|+1}(1-\alpha(t))^{\sum_{l}|Y_{l}|}$
so $p(|Y_{0}|,\dots,|Y_{L}|)$ only depends on $|X_{0}|$ and $M=\sum_{l}|Y_{l}|$. In particular we can sample $M\sim\mathrm{NegativeBinomial}(\alpha(t))$ and then distribute it uniformly into $L+1$ bins.
Proof.
Each $Y_{i}$is generated independently, so we just take the product of probabilities from Prop. F.1. ∎
F.2 Proof of Theorem 4.1
Theorem F.3.
(Proof of Thm. 4.1) Say $X_{0}$ is a sequence with length $L$. Call $q(\cdot\mid L)$ a distribution over sequences of length $L$ which simply samples each letter independently from $\mathrm{Cat}(\pi)$ for a distribution $\pi$ such that $\pi(b)>0$ for all letters $b$. Then, as the number of insertions increases, $M_{1}\to\infty$, $X_{1}$ becomes easier to approximate with $q$:
$\mathrm{KL}(p(X_{1}\mid X_{0},M_{1})||q(X_{1}\mid L+M_{1}))\to 0.$
Proof.
We suppress the subscript $1$. Note by Lem. F.8
$\frac{p(X\mid X_{0},M)}{q(X\mid L+M)}=\frac{\mathrm{ali}(X_{0},X)}{\binom{L+M}{L}\prod_{i=1}^{L}\pi(X_{0}^{(i)})}.$
For a set of $L$ indices $I=i_{1}<i_{2}<\cdots<i_{L}$, call $\chi_{I}=\mathbbm{1}(X_{0}=X^{(i_{1})}\cdots X^{(i_{L})})$. Then $\mathrm{ali}(X_{0},X)=\sum_{I}\chi_{I}$ and $E_{q}\chi_{I}=\prod_{i=1}^{L}\pi(X_{0}^{(i)}).$ Therefore we can write
$E_{p}\log\frac{p(X\mid X_{0},M)}{q(X\mid L+M)}=$ $E_{p}\log\frac{\mathrm{ali}(X_{0},X)}{\binom{L+M}{L}\prod_{i=1}^{L}\pi(X_{0}^{(i)})}$
$=$ $E_{p}\log\frac{\mathrm{ali}(X_{0},X)}{E_{q}\mathrm{ali}(X_{0},X)}$
$\leq$ $E_{p}\left|\frac{\mathrm{ali}(X_{0},X)}{E_{q}\mathrm{ali}(X_{0},X)}-1\right|$
$=$ $\frac{E_{p}\left|\mathrm{ali}(X_{0},X)-E_{q}\mathrm{ali}(X_{0},X)\right|}{E_{q}\mathrm{ali}(X_{0},X)}$
$\leq$ $\frac{E_{p}\left|\mathrm{ali}(X_{0},X)-E_{p}\mathrm{ali}(X_{0},X)\right|}{E_{q}\mathrm{ali}(X_{0},X)}$
$+\frac{\left|E_{p}\mathrm{ali}(X_{0},X)-E_{q}\mathrm{ali}(X_{0},X)\right|}{E_{q}\mathrm{ali}(X_{0},X)}$
$\leq$ $\frac{\mathrm{Std_{p}\left(\mathrm{ali}(X_{0},X)\right)}}{E_{q}\mathrm{ali}(X_{0},X)}+\left|\frac{E_{p}\mathrm{ali}(X_{0},X)}{E_{q}\mathrm{ali}(X_{0},X)}-1\right|.$
We now show that these two terms each go to 0, starting with the second term.
The second term
Say $X$ is generated by picking indices $Z=z_{1}<\dots<z_{L}$ which are $X_{0}$ and then generating all other letters from $\pi$ Say we have indices $I$. Then
$E_{p}\chi_{I}\leq$ $(1-p(I\cap Z=\emptyset))+E_{p}\left[\chi_{I}|I\cap Z=\emptyset\right]$
$=$ $1-\frac{\binom{M+L-L}{L}}{\binom{M+L}{L}}+E_{q}\chi_{I}$
$\leq$ $1-\left(\frac{M-L}{M}\right)^{L}+E_{q}\chi_{I}$
$\leq$ $O\left(L^{2}/M\right)+E_{q}\chi_{I}$
$=$ $(1+o(1))E_{q}\chi_{I}.$
Also
$E_{p}\chi_{I}\geq$ $(1-p(I\cap Z=\emptyset))\times E_{p}\left[\chi_{I}|I\cap Z=\emptyset\right]$
$=$ $\left(1-\frac{\binom{M+L-L}{L}}{\binom{M+L}{L}}\right)\times E_{q}\chi_{I}$
$\geq$ $\left(1-\left(\frac{M}{M+L}\right)^{L}\right)\times E_{q}\chi_{I}$
$\geq$ $\left(1-O\left(L^{2}/M\right)\right)E_{q}\chi_{I}$
$=$ $(1-o(1))E_{q}\chi_{I}.$
Then
$\frac{E_{p}\mathrm{ali}(X_{0},X)}{E_{q}\mathrm{ali}(X_{0},X)}=\frac{\binom{M+L}{L}E_{p}\chi_{I}}{\binom{M+L}{L}E_{q}\chi_{I}}=1+o(1).$
The first term
We first change the expectation in the standard deviation into an expectation over $q$. Say $X$ is generated by picking indices $Z=z_{1}<\dots<z_{L}$ which are $X_{0}$ and then generating all other letters from $\pi$ Say we have indices $I,J$. Then
$E_{p}\chi_{I}\chi_{J}\leq$ $(1-p(I\cap Z,J\cap Z=\emptyset))+E_{p}\left[\chi_{I}\chi_{J}|I\cap Z,J\cap Z=\emptyset\right]$
$\leq$ $1-\frac{\binom{M+L-2\times L}{L}}{\binom{M+L}{L}}+E_{q}\chi_{I}\chi_{J}$
$=$ $O(L^{2}/M)+E_{q}\chi_{I}\chi_{J}.$
We also have from above that
$E_{p}\chi_{I}E_{p}\chi_{J}=(1+o(1))E_{q}\chi_{I}E_{q}\chi_{J}.$
Then
$\mathrm{Var}_{p}\mathrm{ali}(X_{0},X)=\sum_{I,J}\mathrm{Cov}_{p}(\chi_{I},\chi_{J})=\binom{M+L}{L}^{2}o(1)+\sum_{I,J}\mathrm{Cov}_{q}(\chi_{I},\chi_{J}).$
The first term is $o(1)$ against $E_{q}\mathrm{ali}(X_{0},X)^{2}=\binom{M+L}{L}^{2}(E\chi_{I})^{2}$, so we can just focus on the second term, $\mathrm{Var}_{q}\mathrm{ali}(X_{0},X)$.
Note if $I\cap J=\emptyset$ then $\mathrm{Cov}_{q}(\chi_{I},\chi_{J})=0$. Then
$\sum_{J}\mathrm{Cov}_{1}(\chi_{I}\chi_{J})\leq$ $\binom{M+L}{L}-\binom{M+L-L}{L}$
$=$ $o\left(\binom{M+L}{L}\right)$
using the same logic as above. Therefore, $\mathrm{Var}_{q}\mathrm{ali}(X_{0},X)=o\left(\binom{M+L}{L}^{2}\right)=o(E_{q}\mathrm{ali}(X_{0},X)^{2})$. This completes the proof.
F.3 Proof of Theorem 4.2
Theorem F.4.
(Proof of Thm. 4.2) Define $M_{t}$ as the number of mutations up to time $t$, and $\mathrm{prev}(X_{t})$ is the last sequence that gained an insertion to become $X_{t}$. Then the negative log likelihood of a sequence of length $L$, $\log q_{\theta}(X_{0}|L)$, is smaller than
$\mathbb{E}_{M_{t}}\mathrm{KL}$ $(p(X_{1}\mid X_{0},M_{1})||q(X_{1}|L+M_{1}))$
$+\mathbb{E}_{t,X_{t},M_{t}}\frac{M_{t}\beta(t)}{1-\alpha(t)}\mathrm{KL}(p(\mathrm{prev}(X_{t})\mid X_{0},X_{t},M_{t})||q_{\theta}(\mathrm{prev}(X_{t})\mid X_{t},M_{t}))$
Proof.
The proof of Prop. 4.4 from Amin2024-xu derives an ELBO
$\mathbb{E}_{M_{t}}\mathrm{KL}$ $(p(X_{1}\mid X_{0},M_{1})||q(X_{1}|L+M_{1}))$
$+\mathbb{E}_{t,X_{t},M_{t}}w(M_{t},t,X_{0})\mathrm{KL}(p(\mathrm{prev}(X_{t})\mid X_{0},X_{t},M_{t})||q_{\theta}(\mathrm{prev}(X_{t})\mid X_{t},M_{t}))$
where
$w(M_{t},t,X_{0})=\lim_{\epsilon\to 0}E[\#\text{events in }[t-\epsilon,t]|M_{t},X_{0}]/\epsilon.$
The following lemma shows this result. ∎
Lemma F.5.
$w(M,t,X_{0})=M\frac{\beta(t)}{1-\alpha(t)}$
Proof.
First we change out time variable to $\tau=-\log\alpha(t)$. Noting $\log\alpha(t-\epsilon)=\tau-\epsilon\beta(\tau)+O(\epsilon^{2}),$ we have
$w(M,t,X_{0})=\beta(t)\lim_{\epsilon\to 0}\tilde{\mathbb{E}}[\#\text{events in }[\tau-\epsilon,\tau]|M_{\tau}=M]/\epsilon$ (6)
where $\tilde{\mathbb{E}}$ is as if the rate $\beta$ were constant. In SCUD, events occur uniformly in the time interval, so the RHS would be $M/\tau=M/(-\log\alpha(t)).$ For SCISOR, events are more concentrated later in time since more insertions increases the rate of insertion.
$\tilde{\mathbb{E}}[\#\text{events in }[\tau-\epsilon,\tau]|M_{\tau}=M]=$ $P[M_{\tau-\epsilon}=M-1|M_{\tau}=M]+O(\epsilon^{2})$ (7)
$=$ $\frac{P[M_{\tau-\epsilon}=M-1]}{P[M_{\tau}=M]}P[M_{\tau}=M|M_{\tau-\epsilon}=M-1]+O(\epsilon^{2}).$
Note
$P[M_{\tau}=M\mid M_{\tau-\epsilon}=M-1]=$ $P[M_{\tau}\geq M\mid M_{\tau-\epsilon}=M-1]+O(\epsilon^{2})$ (8)
$=$ $P(\mathrm{Exp}(M+|X_{0}|)\leq\epsilon)+O(\epsilon^{2})$
$=$ $1-e^{-\epsilon(M+|X_{0}|)}+O(\epsilon^{2})$
$=$ $\epsilon(M+|X_{0}|)+O(\epsilon^{2}).$
Finally,
$\frac{P[M_{\tau}=M-1]}{P[M_{\tau}=M]}=$ $\frac{\mathrm{NegBin}(|X_{0}|,e^{-\tau};M-1)}{\mathrm{NegBin}(|X_{0}|,e^{-\tau};M)}$ (9)
$=$ $\frac{\binom{m-1+|X_{0}|}{m-1}(1-e^{-\tau})^{M-1}}{\binom{M+|X_{0}|}{M}(1-e^{-\tau})^{M}}$
$=$ $\frac{M}{(M+|X_{0}|)(1-e^{-\tau})}.$
This gives us
$w(M,t,X_{0})=M\frac{\beta(t)}{1-\alpha(t)}$
which is similar for small alpha to the SCUD weight of $w(M,t,X_{0})=M\frac{\beta(t)}{-\log\alpha(t)}$ but becomes larger at larger values. ∎
F.4 Proof of Proposition 4.3
Proposition F.6.
(Proof of Prop. 4.3) Call $\mathrm{ali}(X,Y)$ the number of ways to align a sequence $X$ to a sequence $Y$.
$p(\mathrm{prev}(X_{t})|X_{0},X_{t},M_{t})=\frac{\mathrm{ali}(X_{0},\mathrm{prev}(X_{t}))}{M_{t}\cdot\mathrm{ali}(X_{0},X_{t})}.$
Proof.
Say $Y_{t}$ is $X_{t}$ with a single deletion, the letter $b$. By Lem. F.8
$p(Y_{t}\mid X_{0},X_{t},M_{t})=$ $\frac{p(Y_{t}\mid X_{0},M_{t}-1)}{p(X_{t}\mid X_{0},M_{t})}p(X_{t}\mid Y_{t})$
$=$ $\frac{\binom{L+M_{t}-1}{L}^{-1}\mathrm{ali}(X_{0},Y_{t})}{\binom{L+M_{t}}{L}^{-1}\mathrm{ali}(X_{0},X_{t})\pi(b)}\pi(b)(L+M_{t})^{-1}$
$=$ $\frac{(L+M_{t})\mathrm{ali}(X_{0},Y_{t})}{M_{t}\mathrm{ali}(X_{0},X_{t})}$
$=$ $\frac{\mathrm{ali}(X_{0},Y_{t})}{M_{t}\mathrm{ali}(X_{0},X_{t})}.$
Note finally that we’ve ignored that there may be multiple deletions that take $X_{t}$ to $Y_{t}$ when calculating $p(X_{t}\mid Y_{t})$. We can safely do so as it does not affect the loss Eqn. 3 of any of our other algorithms. ∎
F.5 Derivation of prior matching KL term
Proposition F.7.
(Proof of Prop. A.1) $\mathrm{KL}(p(X_{1}\mid X_{0},M_{1})||q(X_{1}|L+M_{1}))$ is equal to
$\mathbb{E}_{X_{1}\mid X_{0},M_{1}}\left[\log\binom{M_{1}+L}{L}+\sum_{i=1}^{L}\log\pi(X_{0}^{(i)})-\log\mathrm{ali}(X_{0},X_{1})\right].$
Proof.
From Lem. F.8,
$p(X_{1}\mid X_{0},M_{1})=\binom{M_{1}+L}{L}^{-1}\mathrm{ali}(X_{0},X_{1})\prod_{b\in X_{1}\setminus X_{0}}\pi(b).$
Given that $q(X_{1}|L+M_{1})=\prod_{b\in X_{1}}\pi(b)$, this finishes the proof. ∎
F.6 Useful lemma
Lemma F.8.
Calling the letters in $X_{t}$ that are not in $X_{0}$ $X_{t}\setminus X_{0}$,
$p(X_{t}\mid X_{0},M_{t})=\binom{L+M_{t}}{L}^{-1}\mathrm{ali}(X_{0},X_{t})\prod_{b\in X_{t}\setminus X_{0}}\pi(b)$
Proof.
To generate $X_{t}$ from $X_{0},M_{t}$, we could (1) decide which positions $i_{1},\dots,i_{L}\in 1,\dots,L+M_{t}$ should come from $X_{0}$ and then generate the rest of the letters according to $\pi$. Then
$p(X_{t}\mid X_{0},M_{t})=$ $\sum_{\text{indices}\ i_{1},\dots,i_{L}}\frac{\mathbbm{1}(X_{0}=X_{t}^{(i_{1})}\cdots X_{1}^{(i_{l})})}{\binom{L+M_{t}}{L}}\times\prod_{b\in X_{t}\setminus X_{0}}\pi(b)$
$=$ $\binom{L+M_{t}}{L}^{-1}\mathrm{ali}(X_{0},X_{t})\prod_{b\in X_{t}\setminus X_{0}}\pi(b).$
Appendix G Alignments algorithm
Both KL terms in the ELBO make use of the primitive $\textrm{ali}(X,Y)$. In particular, the denoising KL term requires computing the number of alignments between $X_{0}$ and each possible $\mathrm{prev}(X_{t})$, a total of $|X_{t}|$ computations. Naively, computing the alignments between each pair of sequences takes $\mathcal{O}(|X_{0}|\cdot|X_{t}|)$ time for a total of $\mathcal{O}(|X_{0}|\cdot|X_{t}|^{2})$. However, we devise an efficient dynamic programming algorithm to compute all of the alignment terms in parallel in $\mathcal{O}(|X_{0}|\cdot|X_{t}|)$ time, presented in Algorithm 4.
Algorithm 4 Compute $\mathrm{ali}(X_{t}^{-(l)},X_{0})$ for all $l$ in parallel
0: Sequences $X_{0}$ with $|X_{0}|=L$ and $X_{t}$ with $|X_{t}|=N$
1: Set $\text{matching}[i,j]\leftarrow\mathbb{I}(X_{0}^{(i)}=X_{t}^{(j)})$ for all $i\in\{1,\dots,L\}$, $j\in\{1,\dots,N\}$
2: Initialize $\text{prefix\_dp}\leftarrow\mathbf{0}^{(N+1)\times(L+1)}$
3: Set $\text{prefix\_dp}[i,0]\leftarrow 1$ for all $i\in\{1,\dots,N\}$
4:for $l=1$ to $L$ do
5:  $\text{prod}\leftarrow\text{prefix\_dp}[1:N,l-1]\times\text{matching}[l-1,1:N]$
6:  $\text{prefix\_dp}[1:N+1,l]\leftarrow\text{cumsum}(\text{prod},\text{axis}=0)$
7: Initialize $\text{suffix\_dp}\leftarrow\mathbf{1}^{(N+1)\times(L+1)}$
8:for l=L1l=L-1 down to 0 do
9:  $\text{prod}\leftarrow\text{suffix\_dp}[1:N+1,l+1]\times\text{matching}[l,1:N]$
10:  $\text{suffix\_dp}[1:N,l]\leftarrow\text{cumsum}(\text{prod},\text{axis}=0)$
$1]\text{alignments}[l]\leftarrow\sum_{i=1}^{N}\text{prefix\_dp}[i,l]\times\text{suffix\_dp}[i+1,l+1$] for all $l$
12:return alignments